OpenStudy (anonymous):

A line passes through the point (-5,1) and has a slope of -3 . Write an equation for this line.

OpenStudy (anonymous):

\[y-y_1=m(x-x_1)\] m=slope \[point = (x_1,y_1)\]

OpenStudy (afifarahman):

use the formula \[y-y _{1} = m (x-x _{1})\]

OpenStudy (anonymous):

That's how you will start. So it depends what form you want it in If you want it to look like \[y=mx+b\] You will still start with the given equation above.

OpenStudy (anonymous):

ok thanx! ill use this formula